OpenStudy (anonymous):

A challenge - A hikers walks from one town to another. On the 1st day he covers one fifth of the total distance. The next day he covers one quarter of what is left. The following day he covers two fifth of the remainder & on the 4th day, half of what is left. He now has 12 miles left. How many miles has he travelled?

OpenStudy (anonymous):

Start = 100% After Day1 = 80% remaining Day 2 = 60% Day 3 = 36% Day 4 = 18% 12 miles remaining = 18% (12/18)(100) = 66.6 miles I think that's right, correct me if I'm wrong

OpenStudy (dumbcow):

Let total distance be d. day 1: d/5 day2 : (4d/5)*(1/4) = d/5 day3: (3d/5)*(2/5) = 6d/25 day4: (9d/25)*(1/2) = 9d/50 total = 41d/50 Has 9d/50 left 9d/50 = 12 d = 600/9 = 66.66 41d/50 = 54.66 distance traveled is 54.66 miles

OpenStudy (anonymous):

Very close Neophyte, but you're about 11.9 out. it should be 54.7

OpenStudy (anonymous):

Oh yeah I have the total distance, I never subtracted the distance remaining (12) 66.66 - 12 = 54.66 Thanks for that

OpenStudy (anonymous):

yeah that's right..., nice one